Find the volume of the solid S enclosed by the cylinder x2 + y2 = 9 and the planes y + z = 5 and

 = 1. (Hint: Convert to Cylindrical/Polar Coordinates)


1
Expert's answer
2022-01-25T12:29:51-0500
"V=\int \int \int dV=\displaystyle\int_{-3}^{3}\displaystyle\int_{-\sqrt{9-x^2}}^{\sqrt{9-x^2}}\displaystyle\int_{1}^{5-y}dzdydx"

"=\displaystyle\int_{-3}^{3}\displaystyle\int_{-\sqrt{9-x^2}}^{\sqrt{9-x^2}}(4-y)dydx"

"=\displaystyle\int_{0}^{2\pi}\displaystyle\int_{0}^{3}(4-r\sin \theta)rdrd\theta"

"=\displaystyle\int_{0}^{2\pi}[2r^2-\dfrac{r^3}{3}\sin \theta]\begin{matrix}
 3 \\
 0
\end{matrix}d\theta"

"=\displaystyle\int_{0}^{2\pi}(18-9\sin \theta)d\theta"

"=[18\theta+9\cos \theta]\begin{matrix}
 2\pi \\
 0
\end{matrix}=36\pi ({units}^3)"

"36\pi" cubic units.